OpenStudy (anonymous):

Dylan is 9yrs older than Parry. Together their ages sum to 65 yrs. How old is Parry?

OpenStudy (lgbasallote):

let D = dylan's age P= parry's age the sum of their age is 65 D + P = 65 dylan is 9 yrs older than parry so D is 9 yrs older than P D = 9 + P so sub that... D + P = 65 9+P + P = 65 solve for P

OpenStudy (anonymous):

I really do not understand the way you are explanining it to me. so basically what is the answer

OpenStudy (lgbasallote):

are you not understanding the explanation because there is no answer in it?

OpenStudy (anonymous):

no i do not understand the way you are explaining to me

OpenStudy (lgbasallote):

which part?

OpenStudy (anonymous):

all of it

OpenStudy (apoorvk):

@dary seems like we're finding out the answer - only if you could help me with a thing ;) Are you ready?

OpenStudy (anonymous):

sure

OpenStudy (apoorvk):

So, okay. I am Parry and am 'x' years old, and I got a brother Dylan who's is '9' years older to me. So, in terms of 'x', Dylan's age is 9 more than 'x' - agree or not?

OpenStudy (apoorvk):

@dary could I have a bit of your attention please?? :P

OpenStudy (anonymous):

yes sorry

OpenStudy (apoorvk):

so read above and reply

OpenStudy (anonymous):

yes i did

OpenStudy (apoorvk):

I SAY - I am Parry and am 'x' years old, and I got a brother Dylan who's is '9' years older to me. So, in terms of 'x', Dylan's age is 9 more than 'x' - agree or not?

OpenStudy (anonymous):

agree

OpenStudy (apoorvk):

So, 9 more than x is 9 added to x - right?

OpenStudy (apoorvk):

@dary you sleeping or what? just follow me for a minute - okay? do you understand this much that if am 'x' years old and then my brother who is 9 years older to me would be 'x+9' years old?

OpenStudy (anonymous):

ok im listening

OpenStudy (apoorvk):

are you understanding though?? what I just said?

OpenStudy (anonymous):

yes i understand

OpenStudy (apoorvk):

And the sum of their ages that is sum of 'x' and 'x+9', @dary , is 64, so x + x+9 = 64 ----> we have an equation, do you understand this?